OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (0, 0) and (3, 30). Answer 0 5 -10 10

OpenStudy (anonymous):

Use: \[\frac{y_2-y_1}{x_2-x_1}=m\]

OpenStudy (anonymous):

wat?

OpenStudy (anonymous):

Have you not learned the slope formula? Insert x and y points into the formula to solve for m, the slope.

Directrix (directrix):

Slope can be thought of as change in y over change in x. The formula is given above. Continuing, m =[ (30) - 0 ] / [(3 - 0 )] = 30/3 = 10/1. Each time x increases by one, y increases by 10. Slope is 10 ---> answer